Commercial banks in United Economy have total deposits of AED 300 billion. Their reserves are AED 15 billion, two thirds of which are with Central Bank as deposits. There are AED 30 billion notes outside the banks. There are no coins!
Calculate-
a) The monetary base.
b) The banks reserve ratio.
c) Currency drain as a % of deposits
1
Expert's answer
2015-04-30T14:16:58-0400
Deposits of D = AED 300 billion. Reserves are R = AED 15 billion, two thirds of which are with Central Bank as deposits. CU = AED 30 billion notes outside the banks.
a) The monetary base MB = R + CU = 15 + 30 = AED 45 billion.
b) The banks reserve ratio rr = R/D*100% = 15/300*100% = 5%
c) A currency drain is an increase in currency held outside the banks. Currency drain as a % of deposits is 30/300*100% = 10%.